# Through the Looking Glass: Mirror Schrödinger Bridges

### v1

@@ Abstract
Resampling from a target measure whose density is unknown is a fundamental problem in mathematical statistics and machine learning. A setting that dominates the machine learning literature consists of learning a map from an easy-to-sample prior, such as the Gaussian distribution, to a target measure. Under this model, samples from the prior are pushed forward to generate a new sample on the target measure, which is often difficult to sample from directly. In this paper, we propose a new model for conditional resampling called mirror Schrödinger bridges. Our key observation is that solving the Schrödinger bridge problem between a distribution and itself provides a natural way to produce new samples from conditional distributions, giving in-distribution variations of an input data point. We show how to efficiently solve this largely overlooked version of the Schrödinger bridge problem. We prove that our proposed method leads to significant algorithmic simplifications over existing alternatives, in addition to providing control over in-distribution variation. Empirically, we demonstrate how these benefits can be leveraged to produce proximal samples in a number of application domains.
1Massachusetts Institute of Technology   2Columbia University
@@ 1 Introduction
Mapping one probability distribution to another is a central technique in mathematical statistics and machine learning. Myriad computational tools have been proposed for this critical yet often challenging task. Models and techniques for optimal transport provide one class of examples, where methods like the Hungarian algorithm (Kuhn, 1955) map one distribution to another with optimal cost. Adding entropic regularization to the static optimal transport problem yields efficient algorithms like Sinkhorn’s method (Deming & Stephan, 1940; Sinkhorn, 1964), which have been widely adopted in machine learning since their introduction by Cuturi (2013). Static entropy-regularized optimal transportation is equivalent to a dynamical formulation known as the Schrödinger bridge problem (Schrödinger, 1932; Léonard, 2014), which has proven useful to efficiently compute an approximation of the optimal map paired with an interpolant between the input measures.
@@
The most common learning applications of distribution mapping attempt to find a map from a simple prior distribution and a complex data distribution, either using a score-matching strategy (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2021) or leveraging a formulation of the Schrödinger bridge problem (De Bortoli et al., 2021; Shi et al., 2022; 2023; Zhou et al., 2024); other learning applications map one complex data distribution to another (Cuturi, 2013; Courty et al., 2017).
In this paper, we focus instead on the understudied problem of mapping a probability distribution to itself, that is, finding a joint distribution whose marginals are both the same data distribution $\pi$. This task might seem inane at first glance, since two simple couplings satisfy our constraints: one is the independent coupling $p(x,y)=\pi(x)\pi(y)$, and the other is the “diagonal” map given by $p(x,y)=\pi(x)\delta_{y}$. The space of couplings between a measure and itself, however, is far richer than these two extremes and includes models whose conditional distributions are neither identical nor $\delta$ measures.
@@
Despite its simplicity, this setting of the Schrödinger bridge problem suggests a rich application space. Couplings with the same marginal constraints have already proven useful to enhance model accuracy in vision and natural language processing by reinterpreting attention matrices as transport plans (Sander et al., 2022). Few works, however, consider this task from the perspective of optimizing over path measures or provide control over the entropy of the matching at test time. Albergo et al. (2023) propose a stochastic interpolant between a distribution and itself, but their interpolants are not minimal in the relative entropy sense and do not solve the Schrödinger bridge problem, even with optimization. Minimal interpolants in the relative entropy sense are those with minimal kinectic energy, and in applications, minimizing the kinectic energy of a path has been correlated to faster sampling (Shaul et al., 2023).
Contributions. We investigate the mirror Schrödinger bridge problem and demonstrate how it can be leveraged to obtain in-distribution variants of a given input sample. In particular, given a sample $x_{0}\sim p_{\text{data}}$, we build a stochastic process $\{{\mathbf{X}}_{t}\}_{t\in[0,1]}$ with minimal relative entropy under which the sample $x_{0}$ arrives at some $x_{1}\sim p_{\text{data}}$ with $x_{1}$ proximal but not identical to $x_{0}$.
Our contributions in this direction are twofold: first, on the theoretical side, we use the time symmetry of the mirror Schrödinger bridge to prove that it can be obtained as the limit of iterates produced via an alternating minimization procedure; and second, in applications, the implementation of our method allows for sampling from the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$ in such a way that we can control how proximal a generated sample $x_{1}$ is relative to the input sample $x_{0}$.
@@ 2 Related Works
Expectation maximization. Our methodology can be broadly categorized under the umbrella of expectation maximization algorithms, drawing from the theory of information geometry. A number of recent papers introduce related formulations to machine learning; most relevant to us are the works of Brekelmans & Neklyudov (2023); Vargas & Nüsken (2023). These works, however, focus on the case of finding a path measure with two distinct marginal constraints, overlooking the potential application to resampling and algorithmic simplifications obtained for the case in which the marginal constraints are the same. In our work, we derive an algorithm that is distinct, yet similar in flavor, to address this overlooked version of the problem, i.e. the mirror Schrödinger bridge.
Schrödinger bridges and stochastic interpolants. Schrödinger bridges have been used to obtain generative models by flowing samples from a prior distribution to an empirical data distribution from which new data is to be sampled. Several methods have been proposed to this end: De Bortoli et al. (2021); Vargas et al. (2021) iteratively estimate the drift of the SDE associated with the diffusion processes of half-bridge formulations. While the first uses neural networks and score matching, the latter employs Gaussian processes. From these, a number of extensions or alternative methods have been presented; most relevant are (Shi et al., 2023; Peluchetti, 2023), which extend (De Bortoli et al., 2021) but differ with respect to the projection sets used to define their half-bridge formulations.
To the best of our knowledge, the work of Albergo et al. (2023) is the only one in the literature on generative modeling that maps from a distribution to itself. In their paper, flow matching learns a drift function associated with a stochastic path from the data distribution to itself. Their stochastic interpolants, however, are not optimal with respect to any functional. In particular, they lack optimality in the relative entropy sense, a property correlated to sampling effectiveness and generation quality (Shaul et al., 2023) and hence of practical importance. By contrast, our method discovers the coupling with minimal relative entropy, akin to methods such as (De Bortoli et al., 2021; Shi et al., 2023); our method, however, presents certain algorithmic advantages over these, which can only be derived for the mirror case.
@@ 3 Mathematical Preliminaries
Definition. Let $n>0$ be an integer, and let ${\mathbb{P}}^{0}\in\mathcal{P}(C\left([0,1],\mathbb{R}^{n}\right))$ be a reference measure in the space of path measures. Following (Jamison, 1975; Léonard, 2014), we define the Schrödinger bridge problem to be the problem of finding a path measure ${\mathbb{P}}_{\rm{SB}}$ interpolating between prescribed initial and final marginals $\pi_{0}$ and $\pi_{1}$ that is the closest to the reference measure ${\mathbb{P}}^{0}$ with respect to the Kullback-Leibler divergence $D_{\mathrm{KL}}$. To be precise, we define ${\mathbb{P}}_{\rm{SB}}$ to be the solution of the following optimization problem:
@@
The reference path measure ${\mathbb{P}}^{0}$ is typically chosen to be associated with a diffusion process, which is defined to be any stochastic process ${\mathbf{X}}_{t}$ governed by a forward SDE of the form
@@
where $f_{t}$ denotes the forward drift function, $\sigma>0$ is the noise coefficient, and ${\mathbf{W}}_{t}$ denotes the Wiener process. Such a process ${\mathbf{X}}_{t}$ corresponds to a unique path measure once an initial or final condition is specified. An important aspect of diffusion processes is that their time-reversals are diffusion processes of the same noise coefficient $\sigma$. Specifically, if ${\mathbf{X}}_{t}$ is a diffusion process with time-reversal denoted by ${\mathbf{Y}}_{t}$, then ${\mathbf{Y}}_{t}$ is governed by a backward SDE of the form
@@
where $b_{t}$ denotes the backward drift function (see (Winkler et al., 2023, section 2.3)).
In the case where ${\mathbb{P}}^{0}$ arises from a diffusion process, any path measure with finite KL divergence with respect to ${\mathbb{P}}^{0}$, including the Schrödinger bridge ${\mathbb{P}}_{\rm{SB}}$, necessarily also arises from a diffusion process with noise $\sigma$ (Vargas et al., 2021). Consequently, by adjusting the initial condition of the reference SDE, we can assume that the reference process ${\mathbb{P}}^{0}$ has a prescribed initial marginal $\pi_{0}$, without changing the solution to (1).
@@
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\cdot,\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{P}}^{2k}\right),\;{\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi_{0},\cdot)}D_{\mathrm{KL}}\big{(}{\mathbb{P}}\;\|\;{{\mathbb{P}}^{2k+1}}\big{)}$ (2)
where ${\mathbb{D}}(\cdot,\pi_{1})$, respectively, ${\mathbb{D}}(\pi_{0},\cdot)$, denotes the space of path measures with final (resp., initial) marginal fixed to be $\pi_{1}$ (resp., $\pi_{0}$). Ruschendorf (1995) proves that the sequence of iterates ${\mathbb{P}}^{k}$ converges in total variation to ${\mathbb{P}}_{\rm{SB}}$ as $k\to\infty$. IPFP can be thought of as an extension of Sinkhorn’s algorithm to continuous state spaces, where the rescaling updates characteristic of Sinkhorn are replaced by iterated direct $D_{\mathrm{KL}}$ projections onto sets of distributions with fixed initial or final marginal (Essid & Pavon, 2019).
Applications. Suppose $\pi_{0}$ is given by a data distribution $p_{\rm{data}}$ and take $\pi_{1}$ to be an easy-to-sample distribution $p_{\rm{prior}}$, e.g., $p_{\rm{prior}}={\mathcal{N}}(0,{\bm{I}})$. The backward diffusion process associated with ${\mathbb{P}}_{\rm{SB}}$ gives a model for sampling from $p_{\rm{data}}$. In practice, the IPFP iterates in (2) can be solved using an algorithm known as the diffusion Schrödinger bridge (DSB), developed by De Bortoli et al. (2021). DSB relies on the following observation, which is a consequence of Girsanov’s theorem: ${\mathbb{P}}^{2k+1}$ is the path measure whose backward drift is equal to the time-reversal of the forward drift of ${\mathbb{P}}^{2k}$, and ${\mathbb{P}}^{2k+2}$ is the path measure whose forward drift is equal to the time-reversal of the backward drift of ${\mathbb{P}}^{2k+1}$. Leveraging this fact, DSB solves for ${\mathbb{P}}_{\rm{SB}}$ by training neural networks to learn the forward and backward drift functions associated with the IPFP iterates.
@@ 4 Mirror Schrödinger Bridges
Given a reference path measure ${\mathbb{P}}^{0}$ and a prescribed marginal distribution $\pi$, we consider the Schrödinger bridge problem between $\pi$ and itself with respect to ${\mathbb{P}}^{0}$. In the case where ${\mathbb{P}}^{0}$ is time-symmetric, the Schrödinger bridge will inherit the time-symmetry, in which case we call it the mirror Schrödinger bridge from $\pi$ to itself with respect to ${\mathbb{P}}^{0}$. Mathematically, we write
@@
A naïve approach to solving the mirror Schrödinger bridge problem (3) is to apply IPFP with both marginals $\pi_{0}=\pi_{1}$ set equal to $\pi$. In practice, this requires iterative training of two neural networks $f_{t}^{\theta}$ and $b_{t}^{\phi}$, the first modeling the drift of the forward diffusion process associated to ${\mathbb{P}}_{\rm{MSB}}$ and the latter modeling the drift of the corresponding backward process. But this straightforward application of IPFP leads to unnecessary computational expense, as it fails to use the time-symmetry of the problem (3). In particular, at optimality the forward and backward drifts of ${\mathbb{P}}_{\rm{MSB}}$ must be equal, because the mirror Schrödinger bridge ${\mathbb{P}}_{\rm{MSB}}$ is time-symmetric.
@@
In section 4.1, we develop a method for solving (3) by leveraging time-symmetry in conjunction with a general technique from information geometry known as the Alternating Minimization Procedure (AMP), which was first formalized by Csiszár & Tusnády (1984). Then, in section 4.3, we derive an efficient algorithm that involves training a single neural network modeling the drift of the diffusion process associated to ${\mathbb{P}}_{\rm{MSB}}$ and requires half of the computational expense in terms of training iterations for the mirror problem, when compared to other IPFP-based algorithms.
4.1 Alternating Minimization Procedure

### v2

# Mirror Bridges Between Probability Measures

@@ Abstract
Resampling from a target measure whose density is unknown is a fundamental problem in mathematical statistics and machine learning. A setting that dominates the machine learning literature consists of learning a map from an easy-to-sample prior, such as the Gaussian distribution, to a target measure. Under this model, samples from the prior are pushed forward to generate a new sample on the target measure, which is often difficult to sample from directly. Of particular interest is the problem of generating a new sample that is proximate to or otherwise conditioned on a given input sample. In this paper, we propose a new model called mirror bridges to solve this problem of conditional resampling. Our key observation is that solving the Schrödinger bridge problem between a distribution and itself provides a natural way to produce new samples from conditional distributions, giving in-distribution variations of an input data point. We demonstrate how to efficiently estimate the solution to this largely overlooked version of the Schrödinger bridge problem, and we prove that under mild conditions, the difference between our estimate and the true Schrödinger bridge can be controlled explicitly. We show that our proposed method leads to significant algorithmic simplifications over existing alternatives, in addition to providing control over in-distribution variation. Empirically, we demonstrate how these benefits can be leveraged to produce proximal samples in a number of application domains.
@@ 1 Introduction
Mapping one probability distribution to another is a central technique in mathematical statistics and machine learning. Myriad computational tools have been proposed for this critical yet often challenging task. Models and techniques for optimal transport provide one class of examples, where methods like the Hungarian algorithm (Kuhn, 1955) map one distribution to another with optimal cost. Adding entropic regularization to the static optimal transport problem yields efficient algorithms like Sinkhorn’s method (Deming and Stephan, 1940; Sinkhorn, 1964), which have been widely adopted in machine learning since their introduction by Cuturi (2013). Static entropy-regularized optimal transportation is equivalent to a dynamical formulation known as the Schrödinger bridge problem (Schrödinger, 1932; Léonard, 2014), which has proven useful to efficiently compute an approximation of the optimal map paired with an interpolant between the input measures.
@@
The most common learning applications of distribution mapping attempt to find a map from a simple prior distribution to a complex data distribution, either using a score-matching strategy (Song and Ermon, 2019; Ho et al., 2020; Song et al., 2021) or leveraging a formulation of the Schrödinger bridge problem (De Bortoli et al., 2021; Shi et al., 2022, 2023; Zhou et al., 2024); other learning applications map one complex data distribution to another (Cuturi, 2013; Courty et al., 2017).
Rather than mapping a simple prior to a complex data distribution, in this paper we instead tackle the understudied problem of mapping a probability distribution to itself, that is, finding a joint distribution whose marginals are both the same data distribution $\pi$. This task might seem inane at first glance, since two simple couplings satisfy our constraints: one is the independent coupling $p(x,y)=\pi(x)\pi(y)$, and the other is the “diagonal” map given by $p(x,y)=\pi(x)\delta_{y}$. The space of couplings between a measure and itself, however, is far richer than these two extremes and includes models whose conditional distributions are neither identical nor Dirac measures.
Figure 1: We visualize the resampling of a 2D distribution obtained with our method for different values of noise $\sigma$. Higher noise value results in greater in-distribution sample variance.
@@
Despite its simplicity, the mirror case of the Schrödinger bridge problem suggests a rich application space. Couplings with identical marginals have proven useful to enhance model accuracy in vision and natural language processing by reinterpreting attention matrices as transport plans (Sander et al., 2022). Few works, however, consider this task from the perspective of optimizing over path measures or provide control over the entropy of the matching at test time. Albergo et al. (2023) propose a stochastic interpolant between a distribution and itself, but their interpolants are not minimal in the relative entropy sense. Such minimal interpolants are those with minimal kinectic energy, and in applications, minimizing the kinectic energy of a path has been correlated to faster sampling (Shaul et al., 2023).
Contributions. We investigate the mirror Schrödinger bridge problem and demonstrate how it can be leveraged to obtain in-distribution variants of a given input sample. In particular, given a sample $x_{0}\sim p_{\text{data}}$, we build an estimate for the stochastic process $\{{\mathbf{X}}_{t}\}_{t\in[0,1]}$ with minimal relative entropy under which the sample $x_{0}$ arrives at some $x_{1}\sim p_{\text{data}}$ with $x_{1}$ proximal but not identical to $x_{0}$. We call our estimate a mirror bridge. Although this estimate might not be the mirror Schrödinger bridge, we prove that under mild conditions the former is a good approximation of the latter in an explicitly quantifiable way. Furthermore, we demonstrate that our algorithm for obtaining mirror bridges has computational advantages over alternatives used to compute Schrödinger bridges.
Our contributions are twofold. First, on the theoretical side, we use the time symmetry of the mirror Schrödinger bridge to express it as the limit of iterates produced by a symmetrized version of the Iterative Proportional Fitting Procedure (IPFP). We show that this scheme admits a convenient first-order approximation in terms of a convergent alternating minimization procedure that dramatically reduces computational expense in practice. Our mirror bridge is the limit of the latter procedure, and we provide an error bound to justify it as a good approximation. Second, in applications, the implementation of our method allows for sampling from the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$, allowing us to control how proximal a generated sample $x_{1}$ is relative to the input sample $x_{0}$.
@@ 2 Related Works
Expectation maximization. Our method can be categorized under the umbrella of expectation maximization algorithms, drawing from the theory of information geometry. A number of recent papers introduce related formulations to machine learning; most relevant to us are Brekelmans and Neklyudov (2023); Vargas and Nüsken (2023); Vargas et al. (2024). These works, however, focus on finding a path measure with distinct marginal constraints, overlooking the potential application to resampling and algorithmic simplifications obtained for the case in which the marginal constraints are the same.
Schrödinger bridges and stochastic interpolants. Schrödinger bridges have been used to obtain generative models by flowing samples from a prior distribution to an empirical data distribution from which new data is to be sampled. Several methods have been proposed to this end: De Bortoli et al. (2021); Vargas et al. (2021) iteratively estimate the drift of the SDE associated with the diffusion processes of half-bridge formulations. While the first uses neural networks and score matching, the latter employs Gaussian processes. From these, a number of extensions or alternatives have been presented; most relevant are (Shi et al., 2023; Peluchetti, 2023), which extend (De Bortoli et al., 2021) but differ with respect to the projection sets used to define their half-bridge formulations. Schrödinger- bridge-based methods alleviate the computational expense incurred by score-based generative models (SGM) (De Bortoli et al., 2021). The latter requires the forward diffusion process to run for longer times with smaller step sizes. Unlike SGM, our method provides a tool to flow an existing sample in the same data distribution with control over the spread of the newly obtained sample.
To the best of our knowledge, the work of Albergo et al. (2023) is the only one in the literature on generative modeling that considers maps from a distribution to itself. In their paper, flow matching learns a drift function associated with a stochastic path from the data distribution to itself. Their stochastic interpolants, however, do not attempt to seek optimality in the relative entropy sense, a property correlated to sampling effectiveness and generation quality (Shaul et al., 2023). By contrast, our method seeks to estimate the coupling with minimal relative entropy, similar in spirit to methods such as (Vargas et al., 2021; De Bortoli et al., 2021; Shi et al., 2023); our method, however, presents certain algorithmic advantages over these, which can only be derived for the mirror case.
@@ 3 Mathematical Preliminaries
Definition. Let $n>0$ be an integer, and let ${\mathbb{P}}^{0}\in\mathcal{P}(C\left([0,1],\mathbb{R}^{n}\right))$ be a reference measure in the space of path measures. Following (Jamison, 1975; Léonard, 2014), we define the Schrödinger bridge problem as the problem of finding a path measure ${\mathbb{P}}_{\rm{SB}}$ interpolating between prescribed initial and final marginals $\pi_{0}$ and $\pi_{1}$ that is the closest to ${\mathbb{P}}^{0}$ with respect to Kullback-Leibler divergence $D_{\mathrm{KL}}$. To be precise, we define ${\mathbb{P}}_{\rm{SB}}$ to be the solution of the following optimization problem:
@@
The reference path measure ${\mathbb{P}}^{0}$ is typically chosen to be associated with a diffusion process, which is defined to be any stochastic process ${\mathbf{X}}_{t}$ governed by a forward SDE
@@
where $f_{t}$ is the forward drift function, $\sigma>0$ is the noise coefficient, and ${\mathbf{W}}_{t}$ is the Wiener process.Such a process ${\mathbf{X}}_{t}$ corresponds to a unique path measure once an initial or final condition is specified. An important aspect of diffusion processes is that their time-reversals are diffusion processes of the same noise coefficient $\sigma$. That is, if ${\mathbf{X}}_{t}$ is a diffusion process with time-reversal ${\mathbf{Y}}_{t}$, then the ensembles $\left({\mathbf{Y}}_{t}\right)_{t\in[0,1]}$ and $\left({\mathbf{X}}_{1-t}\right)_{t\in[0,1]}$ coincide, and ${\mathbf{Y}}_{t}$ is governed by a backward SDE
@@
where $b_{t}$ denotes the backward drift function (see (Winkler et al., 2023, §2.3)).
In the case where ${\mathbb{P}}^{0}$ arises from a diffusion process, any path measure with finite KL divergence with respect to ${\mathbb{P}}^{0}$, including the Schrödinger bridge ${\mathbb{P}}_{\rm{SB}}$, necessarily also arises from a diffusion process with noise $\sigma$ (Vargas et al., 2021; Léonard, 2014). Consequently, by adjusting the initial condition of the reference SDE, we can assume that the reference process ${\mathbb{P}}^{0}$ has a prescribed initial marginal $\pi_{0}$, without changing the solution to (1).
@@
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\cdot,\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{P}}^{2k}\right),\quad{\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi_{0},\cdot)}D_{\mathrm{KL}}\big{(}{\mathbb{P}}\;\|\;{{\mathbb{P}}^{2k+1}}\big{)},$ (2)
where ${\mathbb{D}}(\cdot,\pi_{1})$ (resp., ${\mathbb{D}}(\pi_{0},\cdot)$) denotes the space of path measures with final (resp., initial) marginal $\pi_{1}$ (resp., $\pi_{0}$). Ruschendorf (1995) proves that the sequence of iterates ${\mathbb{P}}^{k}$ converges in total variation to ${\mathbb{P}}_{\rm{SB}}$ as $k\to\infty$. IPFP can be thought of as an extension of Sinkhorn’s algorithm to continuous state spaces, where the rescaling updates characteristic of Sinkhorn are replaced by iterated direct $D_{\mathrm{KL}}$ projections onto sets of distributions with fixed initial or final marginal (Essid and Pavon, 2019).
Applications. Suppose $\pi_{0}$ is given by a data distribution $p_{\rm{data}}$, and take $\pi_{1}$ to be an easy-to-sample distribution $p_{\rm{prior}}$, e.g., ${\mathcal{N}}(0,{\bm{I}})$. The backward diffusion process associated with ${\mathbb{P}}_{\rm{SB}}$ gives a model for sampling from $p_{\rm{data}}$. In practice, the IPFP iterates in (2) can be solved using diffusion Schrödinger bridge (DSB), an algorithm developed by De Bortoli et al. (2021). DSB relies on the following fact, which is a consequence of Girsanov’s theorem: ${\mathbb{P}}^{2k+1}$ is the path measure whose backward drift equals the time-reversal of the forward drift of ${\mathbb{P}}^{2k}$, and ${\mathbb{P}}^{2k+2}$ is the path measure whose forward drift equals the time-reversal of the backward drift of ${\mathbb{P}}^{2k+1}$. Leveraging this, DSB solves for ${\mathbb{P}}_{\rm{SB}}$ by training neural networks to learn the forward and backward drifts associated with the IPFP iterates.
@@ 4 Mirror Schrödinger Bridges
Given a reference path measure ${\mathbb{P}}^{0}$ and a prescribed marginal distribution $\pi$, we consider the Schrödinger bridge problem between $\pi$ and itself with respect to ${\mathbb{P}}^{0}$. In the case where ${\mathbb{P}}^{0}$ is time-symmetric, meaning that the path measure is invariant under the transformation $t\to-t$, the Schrödinger bridge will inherit the time-symmetry
@@
A naïve approach to solving the mirror Schrödinger bridge problem (3) is to apply IPFP with both marginals $\pi_{0}=\pi_{1}$ set equal to $\pi$. In practice, this requires iterative training of two neural networks $f_{t}^{\theta}$ and $b_{t}^{\phi}$, the first modeling the drift of the forward diffusion process associated to ${\mathbb{P}}_{\rm{MSB}}$ and the latter modeling the drift of the corresponding backward process. But this straightforward application of IPFP leads to unnecessary computational expense, as it fails to use the time-symmetry of the problem (3). In particular, at optimality the forward and backward drifts of ${\mathbb{P}}_{\rm{MSB}}$ must be equal, because the mirror Schrödinger bridge ${\mathbb{P}}_{\rm{MSB}}$ is time-symmetric. Related works in entropic optimal transportation suggest that the use of one optimization variable for the static transport formulation in the symmetric case (see (Kurras, 2015, Section 3) and (Feydy et al., 2019, Equations (24)-(25))), but to our knowledge no approach has been developed to leverage symmetry for the dynamical formulation in the language of path measures. This suggests the strategy of modeling the drift with a single neural network and using the IPFP iterations to recursively train that network. While this strategy already yields a significant simplification in the implementation of IPFP, it turns out that we can do even better by further exploiting the time-symmetry of the mirror problem.
@@
In §4.1, we introduce a symmetrized version of IPFP, replacing the second projection with a step that enforces symmetry of the bridge iterate. In §4.2, we present the key insight that, at least to first order, this symmetrization step can be performed analytically by averaging the forward and backward drifts of the previous iterate. We show that this drift averaging can be interpreted as a minimization problem in its own right, by simply reversing the direction of the $D_{\mathrm{KL}}$ divergence in the symmetrization step. The resulting method is a convenient first-order approximation of symmetrized IPFP in terms of what is known in information geometry as the Alternating Minimization Procedure (AMP), first formalized by Csiszár and Tusnády (1984). Then, in §4.3, we derive an efficient algorithm that trains a single neural network modeling the drift associated to ${\mathbb{P}}_{\rm{MSB}}$ and requires half the computational expense in terms of training iterations for the mirror problem compared to other IPFP-based algorithms.
4.1 Symmetrized Iterative Proportional Fitting